OpenStudy (anonymous):

How do I solve for x log3x + log34 = 2

OpenStudy (anonymous):

@aum this is the equation, I have no clue how to do this

OpenStudy (anonymous):

Hint: log39=2

OpenStudy (anonymous):

how'd you get that @Saeeddiscover

OpenStudy (anonymous):

logab=c --> b=a^c

OpenStudy (aum):

\( \Large \log_3x + \log_34 = 2 \\ \Large \log_3(4*x) = 2 \\ \Large 4x = 3^2 = 9 \\ \Large x = 9/4 = 2\frac 14 \)

OpenStudy (aum):

\(\Large \log(AB) = \log(A) + \log(B) \) \(\Large \log_bA = C ~~\text{ implies }~~A = b^C\)

OpenStudy (anonymous):

is that complete ? like step by step

OpenStudy (anonymous):

I'm confused @ aum

OpenStudy (anonymous):

@aum im confused

OpenStudy (jdoe0001):

have you covered the logarithm rules yet? http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif

OpenStudy (phi):

Here is one of the rules you need to use to solve your problem http://www.khanacademy.org/math/algebra2/logarithms-tutorial/logarithm_properties/v/sum-of-logarithms-with-same-base